# Labels or Input? Rethinking Augmentation in Multimodal Hate Detection

### v1

@@ Abstract
The modern web is saturated with multimodal content, intensifying the challenge of detecting hateful memes, where harmful intent is often conveyed through subtle interactions between text and image under the guise of humor or satire. While recent advances in Vision-Language Models (VLMs) show promise, these models lack support for fine-grained supervision and remain susceptible to implicit hate speech. In this paper, we present a dual-pronged approach to improve multimodal hate detection. First, we propose a prompt optimization framework that systematically varies prompt structure, supervision granularity, and training modality. We show that prompt design and label scaling both influence performance, with structured prompts improving robustness even in small models, and InternVL2 achieving the best F1-scores across binary and scaled settings. Second, we introduce a multimodal data augmentation pipeline that generates 2,479 counterfactually neutral memes by isolating and rewriting the hateful modality. This pipeline, powered by a multi-agent LLM–VLM setup, successfully reduces spurious correlations and improves classifier generalization. Our approaches inspire new directions for building synthetic data to train robust and fair vision-language models. Our findings demonstrate that prompt structure and data composition are as critical as model size, and that targeted augmentation can support more trustworthy and context-sensitive hate detection.
@@
Datasetshttps://hf.co/datasets/sahajps/Meme-Sanity
@@ Introduction
The rise of Vision-Language Models (VLMs) such as CLIP (Radford et al. 2021), Flamingo (Alayrac et al. 2022), and InternVL2 (Chen et al. 2024) has led to significant progress in understanding and processing multimodal content. These models are increasingly applied to downstream tasks, including meme classification and online content moderation. At the same time, visual harmful content has surged across platforms like X and Instagram, where images and videos constitute between 30% of shared content on text-centric platforms and nearly 100% on image-first platforms (Peng, Lu, and Shen 2023; Yang, Davis, and Hindman 2023; Pfeffer et al. 2023). This shift underscores the growing need for scalable, robust methods that can effectively address the complexities of multimodal hate speech detection in contemporary social media environments (González-Aguilar, Segado-Boj, and Makhortykh 2023; Heley, Gaysynsky, and King 2022; Solea and Sugiura 2023). Moreover, recent studies find that VLMs are more susceptible to harmful content than their unimodal counterparts, with image-text combinations inducing unsafe behaviors even under benign conditions (Hee and Lee 2025).
@@
While prior work has addressed aspects of hate meme detection, many existing approaches focus on fine-tuning pre-trained models without exploring how prompting might be used to enrich supervision or guide classification (Lippe et al. 2020; Hermida and Santos 2023). Although prompt engineering has shown promise in unimodal tasks as a lightweight method to influence model behavior (Furniturewala et al. 2024), its utility in the multimodal setting, particularly for enhancing training labels or shaping classification decisions in VLMs, remains underexplored. These research gaps motivate our core research question: Is improving prompting or refining the underlying data more effective than a standard baseline in multimodal hateful content classification?
@@
Figure 1: Overview of the augmentation pipeline. Red and green boxes indicate hateful and non-hateful content. In part B, the augmented Data is created for the green-red pairs – cases where the background is non-hateful but is used in combination with a hateful caption, resulting in a hateful interpretation.
@@
Recent studies also highlight significant limitations in dataset quality and data augmentation practices. Several datasets are constructed through manual pipelines that involve overlaying text on stock images, yielding templated meme formats (Nandi, Sharma, and Chakraborty 2024). Broader surveys on multimodal hate content (Hee et al. 2024) emphasize the growing presence of hateful speech across text, image, and video, but often stop short of proposing actionable modeling strategies. Synthetic augmentation methods such as MixGen (Hao et al. 2023) attempt to increase diversity by randomly mixing images and captions, but frequently produce semantically incoherent pairings that reduce realism. In parallel, cultural and representational biases, and stereotypes embedded in VLMs (Madasu, Lal, and Howard 2025; Liu et al. 2025) are increasingly recognized yet get very little attention in multimodal hate speech detection, especially culturally grounded hateful memes (Bui, von der Wense, and Lauscher 2025).
@@
In this study, we investigate two complementary strategies (see Figure 1) for improving hateful meme classification: (1) prompt optimization, which leverages structured prompting to enrich supervision across modalities and label formats; and (2) multimodal augmentation, which generates counterfactually neutral memes to reduce spurious correlations.
@@
We conduct experiments on the Facebook Hateful Memes dataset (Kiela et al. 2020), which contains over 10K image-caption pairs labeled for binary hatefulness. Our evaluation involves fine-tuning both unimodal and multimodal models under each strategy. This setup enables a direct comparison of prompt-informed supervision and data-centric augmentation with standard baselines for robust multimodal classification. Our study finds that label augmentation, such as scaling hatefulness, shows consistent patterns across modalities and prompting styles. This indicates that well-augmented multimodal hate labels can enable a more nuanced understanding of harmful content beyond simple binary classification. We also observe that input-level multimodal augmentation improves performance across modalities when fine-tuning pretrained models. Human evaluation further supports the quality of the augmented data, highlighting its potential in reducing toxicity and pointing to the broader promise of large multimodal models when used strategically in multi-agent systems.
Contributions.
@@
  • We extend the Facebook Hateful Memes dataset by generating 2479 counterfactually neutral memes, offering a novel benchmark for evaluating robustness to spurious text-image correlations.
  • We introduce a structured prompt optimization framework for multimodal classification, demonstrating that prompt design—long studied in language-only settings—has overlooked yet significant effects in vision-language tasks.
@@
Prompting as a mechanism for supervision.
@@
Augmentation as a tool for improving robustness.
@@
For our first set of experiments, we explore a full-factorial design spanning 3 learning styles (text-only: fine-tuning, multimodal: fine-tuning & prompting), 2 prompt strategies (simple vs. category-based), and 2 output types (binary vs. scale-based). We use prompting as a labeling instrument to generate enriched supervision signals. Specifically, we design structured prompt templates - ranging from simple to category-specific prompts - and from binary to scale-based outputs, and use them with a teacher model (GPT-4o-mini [^1] https://platform.openai.com/docs/models/gpt-4o-mini) on a subset of training data (due to high cost associated with the use of GPT-4o-mini with vision inputs) to support scale-based labeling and generate interval annotations (0–9), thereby enhancing label granularity. These enhanced labels are then used to fine-tune both a large multimodal model (InternVL2 8B (Chen et al. 2024)) and a smaller unimodal model (DistilBERT (Sanh et al. 2020)), allowing us to evaluate the impact of prompt-informed supervision. All prompt variations are carefully controlled and documented to ensure consistency (refer to supplementary information (SI)).
@@
All prompt templates are modular and composable with both label formats (see details in SI). We fix prompt wording and structure across conditions to minimize confounds.
@@
To enable fine-tuning on scale outputs, we use a teacher model to generate soft labels for training. Incorrect labels (i.e., those inconsistent with the binary ground truth) are manually filtered out to ensure label quality.
@@
For our second set of experiments, we explore a data-centric augmentation approach, applying a multi-step and multi-agent pipeline to regenerate non-hateful variants of memes using LLM and VLM tools. This approach produces a bias-aware augmented dataset designed to reduce spurious correlations while preserving semantic and visual coherence.
@@
This augmentation targets memes for which the caption is the primary hateful modality—i.e., where $I_{i}$ is semantically benign. The hypothesis is that injecting such counterfactuals will decorrelate textual toxicity from visual features and improve generalization.
@@
Full implementation details, including model prompts and architecture-specific settings, are provided in the SI.
@@
We now describe the datasets, baselines, and models used to evaluate our two experimental approaches, prompt optimization and multimodal augmentation, for the task of classifying hate memes. Our evaluation focuses on classification accuracy and weighted-F1 score, as standard in the Facebook Hateful Memes benchmark. Note that in each setting, we are splitting the training set into two parts: one for training and one for validation. For testing, we are combining both seen and unseen test data. Additional details, including prompt templates and training configurations, are provided in the SI.
@@
We use the Facebook Hateful Memes dataset (Kiela et al. 2020), a benchmark corpus for multimodal hate speech detection. The dataset consists of over 10,000 image-caption pairs for training and 3,000 for test evaluation. Each meme is annotated with a binary label indicating whether it is hateful. Captions are written in natural language and paired with images that range from benign stock photos to visually suggestive content.
For the prompt optimization experiments, we use the original dataset labels (binary) and augment them with scale-based annotations using GPT-4o-mini as a teacher model.
For the multimodal augmentation experiments, we extract a subset of the training set in which hatefulness is localized to the caption only (i.e., image visuals are neutral). These samples are regenerated using our multi-step and multi-agent pipeline to create counterfactually non-hateful memes while preserving visual context.
@@
To evaluate both label augmentation and multimodal data augmentation strategies, we fine-tune (sometimes prompt, where applicable, such as InternVL2) a suite of unimodal and multimodal models on the Facebook Hateful Memes dataset. For fine-tuning, AdamW is used as an optimizer with cross-entropy as a loss function, which is standard in classification tasks.
@@
We compare performance across zero-shot prompting with InternVL2 and prompt-informed fine-tuning, using InternVL2 and DistilBERT. All smaller models are fully fine-tuned with the same hyperparameters to ensure comparability. For InternVL2, we apply LoRA (Hu et al. 2022) -based parameter-efficient tuning. The learning rate and batch size are taken as $10^{-4}$ and $16$ respectively. Results and model sizes are given in Table 2.
@@
For this augmentation track, we fine-tune each pre-trained model on both the original dataset and an extended version $\tilde{D}$, which includes counterfactually neutral memes. We experiment with three model types: unimodal (text), unimodal (vision), and multimodal (text + vision), allowing a direct comparison of robustness to spurious correlations. To address class imbalance, we apply downsampling. Fine-tuning is performed on the last two layers of the encoders, with a learning rate of $2\times 10^{-5}$ and batch size of 128. Table 3 presents the accuracy and F1-scores for all models.
Models used are publicly available checkpoints or APIs. Hyperparameters are also kept consistent across settings to control for confounding effects. Full training and inference configurations are included in the SI.
@@
Original (Binary) Scaled
@@
Unimodal
(fine-tuning)
[…]
Multimodal
[…]
Multimodal
(fine-tuning)
[…]
Table 2: Classification results with binary and scaled labels on simple and category styled prompts. The original dataset has the binary labels. Scaled labels were granular values generated by a teacher model, GPT-4o-mini.
@@
Original (Binary) Extended
@@
Table 3: Classification results with and without newly augmented non-hateful meme data. Note: Until specified, we have used the base variant of the models mentioned. Note that the size is written as an approximate number of parameters in the model.
@@
We present the results and implications of our data augmentation strategy using binary and scaled labels, evaluated via prompting and fine-tuning. While we report both accuracy and weighted F1-score, we focus on F1-score due to its greater sensitivity in class-imbalanced classification tasks.
@@
Under the binary label setting, the category prompt improves performance for multimodal prompting with InternVL2 (F1: 63.00 vs. 61.36), and further boosts it under fine-tuning (F1: 68.30 vs. 67.51). For scaled labels, results are mixed: a slight decrease with prompting and a minor increase (F1: 64.74 vs. 64.63) with fine-tuning. Since the scaled dataset is a small subset of the full training data, these improvements may grow with larger datasets.
Surprisingly, for DistilBERT, the category prompt underperforms the simple prompt in both cases (F1: 49.46 vs. 59.87 and 55.58 vs. 58.41), likely due to overloading the smaller model with excessive category information.
@@
Multimodal fine-tuning with InternVL2 achieves the highest overall F1-scores: 68.30 (binary) and 64.74 (scaled), outperforming prompting-only setups across the board. This highlights the advantage of full model adaptation for downstream hatefulness classification. Even so, zero-shot prompting with InternVL2 yields competitive performance (F1: 63.00 binary; 63.89 scaled), indicating strong generalization capabilities with appropriate prompts.
@@
  • Prompt quality matters: Structured, informative prompts (category) improve performance in most cases, particularly for larger models.
  • Scaled labels are learnable: Despite added complexity, models trained on scaled annotations achieve robust performance (e.g., 64.74 F1 for InternVL2), enabling nuanced hatefulness detection.
  • Lightweight deployment is viable: Even DistilBERT achieves 55.58 F1 on scaled labels with category prompts, showing promise for low-resource or real-time applications.
@@
To assess the value of augmenting our training set with counterfactually neutral memes, we fine-tune a diverse range of popular unimodal and multimodal models, including BERT (Devlin et al. 2019), RoBERTa (Liu et al. 2019), ALBERT (Lan et al. 2020), ViT (Wu et al. 2020), BEiT (Bao et al. 2022), DeiT (Touvron et al. 2021), and CLIP (Radford et al. 2021), using identical hyperparameters across both the original and extended datasets. We aim to determine whether exposure to visually and lexically similar non-hateful memes improves model generalization and robustness.
@@
Across most model types, the extended dataset yields modest gains in accuracy (typically <<1%), but more notable improvements in F1-score, indicating better handling of class imbalance and edge cases. For example:
  • BERT improves from 55.50 to 57.95 in F1 (+2.45), with a negligible drop in accuracy (59.08 to 58.20).
  • RoBERTa shows a larger relative F1 gain, from 47.18 to 52.54 (+5.36), despite a slight accuracy decrease.
  • CLIP, the sole multimodal model, is evaluated, improving from 47.71 to 51.77 in F1 (+4.06), showing that the augmentation is also effective in the joint vision-language space.
The only unimodal model that sees near-constant performance is ALBERT, which improves slightly in F1 from 54.96 to 58.54 (+3.58), but with no meaningful change in accuracy. Among vision-only models, BEiT shows the best improvement (F1: 53.38 to 55.43), while DeiT lags with marginal change.
Text Dominance and Modality Gaps
As expected, text-based models outperform vision-only models across both datasets. For instance, on the extended data, BERT and RoBERTa achieve higher F1-scores (57.95 and 52.54, respectively) compared to ViT and DeiT (53.58 and 51.99). This reinforces prior findings (Aggarwal et al. 2024) that the primary signal of hatefulness often resides in the caption, not the image.
These findings corroborate that counterfactual data augmentation enhances model robustness, especially in F1, by reducing overfitting to superficial features and encouraging deeper semantic reasoning. Such augmentation strategies are particularly effective when models have sufficient capacity to leverage subtle distinctions between hateful and non-hateful content, in larger contextual models like RoBERTa and CLIP.
@@
To assess the reliability of the label augmentation strategy, we conduct a human evaluation on a randomly selected sample of 200 memes. Although the newly scaled labels are checked against associated hate labels and passed through a post-filtering step, we ask human annotators to independently indicate whether they agree or disagree with each assigned label. The agreement rate comes out to be 89%, indicating high consistency between human judgment and the augmented labels.

### v2

@@ Abstract
Online hate remains a significant societal challenge, especially as multimodal content enables subtle, culturally grounded, and implicit forms of harm. Hateful memes embed hostility through text–image interactions and humor, making them difficult for automated systems to interpret. Although recent Vision-Language Models (VLMs) perform well on explicit cases, their deployment is limited by high inference costs and persistent failures on nuanced content. This work examines how far small models can be improved through prompt optimization, fine-tuning, and automated data augmentation. We introduce an end-to-end pipeline that varies prompt structure, label granularity, and training modality, showing that structured prompts and scaled supervision significantly strengthen compact VLMs. We also develop a multimodal augmentation framework that generates counterfactually neutral memes via a coordinated LLM–VLM setup, reducing spurious correlations and improving the detection of implicit hate. Ablation studies quantify the contribution of each component, demonstrating that prompt design, granular labels, and targeted augmentation collectively narrow the gap between small and large models. The results offer a practical path toward more robust and deployable multimodal hate-detection systems without relying on costly large-model inference.
footnotetext: Warning: This paper includes instances of hateful memes and offensive language. These don’t reflect the authors’ views.
@@
Datasethttps://hf.co/datasets/sahajps/Meme-Sanity
@@ Introduction
The rise of Vision-Language Models (VLMs) such as CLIP (Radford et al. 2021), Flamingo (Alayrac et al. 2022), and InternVL2 (Chen et al. 2024) has led to significant progress in understanding and processing multimodal content. These models are increasingly applied to downstream tasks, including meme classification and online content moderation. At the same time, visual harmful content has surged across platforms like X and Instagram, where images and videos constitute between 30% of shared content on text-centric platforms and nearly 100% on image-first platforms (Peng, Lu, and Shen 2023; Yang, Davis, and Hindman 2023; Pfeffer et al. 2023). This shift underscores the growing need for scalable, robust methods that can effectively address the complexities of multimodal hate speech detection in contemporary social media environments (González-Aguilar, Segado-Boj, and Makhortykh 2023; Heley, Gaysynsky, and King 2022; Solea and Sugiura 2023). Moreover, recent studies find that VLMs are more susceptible to harmful content than their unimodal counterparts, with image–text combinations inducing unsafe behaviors even under benign conditions (Hee and Lee 2025).
A central concern in computational social science is the quality of data used to study online harm. High-quality data depends on high-quality annotation, yet annotation cannot be carried out arbitrarily: inconsistent or weak supervision introduces artifacts, reinforces cultural biases, and destabilizes moderation systems (Lee et al. 2024; Masud et al. 2024). This creates a methodological gap, particularly acute for multimodal content, between the complexity of online harm and the limited supervision that existing datasets can reliably provide.
@@
While prior work has addressed aspects of hate meme detection, many existing approaches focus on fine-tuning pre-trained models without exploring how prompting might be used to enrich supervision or guide classification (Lippe et al. 2020; Hermida and Santos 2023). Although prompt engineering has shown promise in unimodal tasks as a lightweight method for influencing model behavior (Furniturewala et al. 2024), its utility in multimodal settings, especially for generating higher-quality labels or shaping classification decisions in VLMs, remains underexplored. These research gaps motivate our core research question: Is improving prompting or refining the underlying data more effective than a standard baseline in multimodal hateful content classification?
@@
Figure 1: Overview of the augmentation pipeline. Red and green boxes indicate hateful and non-hateful content. In part B, the augmented data is created for the green–red pairs—cases where the background is non-hateful but is used in combination with a hateful caption, resulting in a hateful interpretation.
@@
Recent studies also highlight significant limitations in dataset quality and data augmentation practices. Several datasets are constructed through manual pipelines that involve overlaying text on stock images, yielding templated meme formats (Nandi, Sharma, and Chakraborty 2024). Broader surveys on multimodal hate content (Hee et al. 2024) emphasize the growing presence of hateful speech across text, image, and video, but often stop short of proposing actionable modeling strategies. Synthetic augmentation methods such as MixGen (Hao et al. 2023) attempt to increase diversity by randomly mixing images and captions, but frequently produce semantically incoherent pairings that reduce realism. In parallel, cultural and representational biases embedded in VLMs (Madasu, Lal, and Howard 2025; Liu et al. 2025) are increasingly recognized yet receive limited attention in multimodal hate speech detection, particularly for culturally grounded hateful memes (Bui, von der Wense, and Lauscher 2025).
@@
In this study, we investigate two complementary strategies (see Figure 1) for improving hateful meme classification: (1) prompt optimization, which leverages structured prompting to enrich supervision across modalities and label formats; and (2) multimodal augmentation, which generates counterfactually neutral memes to reduce spurious correlations. These approaches offer alternatives to enhance dataset quality, mitigate bias, and conduct a thorough multimodal analysis of social media content.
@@
We conduct experiments on the Facebook Hateful Memes dataset (Kiela et al. 2020), which contains over 10K image–caption pairs labeled for binary hatefulness. Our evaluation involves fine-tuning both unimodal and multimodal models under each strategy. This setup enables a direct comparison of prompt-informed supervision and data-centric augmentation with standard baselines for robust multimodal classification. Our study finds that label augmentation, such as scaling hatefulness, shows consistent patterns across modalities and prompting styles. This indicates that well-augmented multimodal hate labels can enable a more nuanced understanding of harmful content beyond simple binary classification. We also observe that input-level multimodal augmentation improves performance across modalities when fine-tuning pretrained models. Human evaluation further supports the quality of the augmented data, highlighting its potential in reducing toxicity and pointing to the broader promise of large multimodal models when used strategically in multi-agent systems.
Contributions
@@
  • We extend the Facebook Hateful Memes dataset by generating 2479 counterfactually neutral memes, offering a novel benchmark for evaluating robustness to spurious text–image correlations.
  • We introduce a structured prompt optimization framework for multimodal classification, demonstrating that prompt design, long studied in language-only settings, has overlooked yet significant effects in vision–language tasks.
@@
Prompting as a mechanism for supervision
@@
Augmentation as a tool for improving robustness
@@
For our first set of experiments, we explore a full-factorial design spanning 3 learning styles (text-only: fine-tuning, multimodal: fine-tuning & prompting), 2 prompt strategies (simple vs. category-based), and 2 output types (binary vs. scale-based). We use prompting as a labeling instrument to generate enriched supervision signals. Specifically, we design structured prompt templates ranging from simple to category-specific prompts and from binary to scale-based outputs, and use them with a teacher model (GPT-4o-mini [^0] https://platform.openai.com/docs/models/gpt-4o-mini) on training data to support scale-based labeling and generate interval annotations (0–9), thereby enhancing label granularity. These enhanced labels are then used to fine-tune both a large multimodal model (InternVL2 4B (Chen et al. 2024)) and a smaller unimodal model (BERT (Devlin et al. 2019), RoBERTa (Liu et al. 2019)), allowing us to evaluate the impact of prompt-informed supervision.
@@
All prompt templates are modular and composable with both label formats (see details in Appendix). We fix prompt wording and structure across conditions to minimize confounds.
@@
To enable fine-tuning on scale outputs, we use a teacher model to generate soft labels for training. Incorrect labels (i.e., those inconsistent with the binary ground truth) are filtered out to ensure label quality.
@@
The Facebook Hateful Memes dataset exhibits several known artifacts, including templated image-text combinations, culturally skewed representations, and a heavy concentration of caption-driven hatefulness. These are the structural limitations that motivate our second set of experiments, involving counterfactual augmentation strategy.
We developed a multi-step and multi-agent pipeline to regenerate non-hateful variants of memes using LLM and VLM tools. This approach produces a bias-aware augmented dataset designed to reduce spurious correlations while preserving semantic and visual coherence.
@@
This augmentation targets memes for which the caption is the primary hateful modality, i.e., where $I_{i}$ is semantically benign. The hypothesis is that injecting such counterfactuals will decorrelate textual toxicity from visual features and improve generalization.
@@
Full implementation details, including model prompts and architecture-specific settings, are provided in the Appendix.
@@
We now describe the datasets, baselines, and models used to evaluate our two experimental approaches, prompt optimization and multimodal augmentation, for the task of classifying hate memes. Our evaluation focuses on classification accuracy and weighted-F1 score, as standard in the hate classification benchmarks.
Example 1 Example 2 Example 3 Example 4
Examples
HS*
9 9 0 0
Table 1: Consistent examples of the scaled labels.
@@
We use the Facebook Hateful Memes dataset (Kiela et al. 2020), a benchmark corpus for multimodal hate speech detection. The dataset consists of 12.9K image-caption pairs for training, 1.04K for validation, and 3K for test evaluation. Each meme is annotated with a binary label indicating whether it is hateful. Captions are written in natural language and paired with images that range from benign stock photos to visually suggestive content.
For the prompt optimization experiments, we use the original dataset labels (binary) and augment the training and validation sets with scale-based annotations, followed by misaligned example removal. By doing this, we end up with approximately 12K and 1K examples for training and validation, respectively.
For the multimodal augmentation experiments, we extract a subset of the training set in which hatefulness is localized to the caption only (i.e., image visuals are neutral). These samples are regenerated using our multi-step and multi-agent pipeline to create counterfactually non-hateful memes while preserving visual context. Using our multi-agent pipeline, we generate 2,479 counterfactually neutral variants, creating an extended dataset. Unlike the prompt optimization track, we do not use the original validation split. Instead, both configurations are balanced via random downsampling and split 80:20 into new train-validation sets, yielding approximately 9.56K/2.39K (Original) and 11.06K/2.77K (Extended) for training/validation, respectively. This ensures controlled distributions in both configurations while allowing an extended set to benefit from additional augmented data.
@@
To evaluate both label augmentation and multimodal data augmentation strategies, we fine-tune (sometimes prompt, where applicable, such as InternVL2) a suite of unimodal and multimodal models on the Facebook Hateful Memes dataset. For fine-tuning, AdamW is used as an optimizer with cross-entropy as a loss function, which is standard in classification tasks. All training was conducted on commodity GPUs (8×NVIDIA H100 GPU cluster) using consistent hyperparameters across models. Models used are publicly available checkpoints or APIs. Hyperparameters are also kept consistent across settings to control for confounding effects. Training and inference configurations are included in the Appendix.
@@
We compare performance across zero-shot prompting with InternVL2 and prompt-informed fine-tuning, using InternVL2 and BERT-based models (BERT, RoBERTa). All smaller models are fully fine-tuned with the same hyperparameters to ensure comparability. For InternVL2, we apply LoRA (Hu et al. 2022)-based parameter-efficient tuning. Results and model sizes are given in Table 3.
@@
For this augmentation track, we fine-tune each pre-trained model on both the original dataset and an extended version, which includes counterfactually neutral memes. We experiment with three types of encoder-based models: unimodal (text), unimodal (vision), and multimodal (text + vision), allowing a direct comparison of robustness to spurious correlations. We also test the current state-of-the-art small VLMs in zero-shot and fine-tuning settings. For encoder-based models, fine-tuning is performed on the last two layers of the model, and for VLMs’ fine-tuning, we apply QLoRA (Dettmers et al. 2023)-based parameter-efficient tuning. Tables 4 and 5 present the accuracy and F1-scores for all models.
@@
Original (Binary) Augmented (Scaled)
@@
Unimodal (fine-tuning) BERT 112M Simple 60.53 $\pm$ 1.78 58.88 $\pm$ 1.87 61.01 $\pm$ 1.70 57.30 $\pm$ 1.86
Category 54.11 $\pm$ 1.99 53.41 $\pm$ 2.08 60.75 $\pm$ 1.85 59.18 $\pm$ 1.96
RoBERTa 125M Simple 61.54 $\pm$ 1.75 59.13 $\pm$ 1.86 60.79 $\pm$ 1.95 57.42 $\pm$ 2.08
Category 53.55 $\pm$ 1.91 52.87 $\pm$ 2.01 61.49 $\pm$ 1.85 58.24 $\pm$ 2.09
Multimodal (prompting) InternVL2 4B Simple 61.63 $\pm$ 1.85 54.51 $\pm$ 2.33 63.49 $\pm$ 2.02 60.00 $\pm$ 2.25
Category 61.81 $\pm$ 1.93 53.11 $\pm$ 2.43 64.46 $\pm$ 1.91 64.74 $\pm$ 1.89
Multimodal (fine-tuning) InternVL2 4B Simple 68.15 $\pm$ 1.77 67.28 $\pm$ 1.83 64.98 $\pm$ 1.79 63.06 $\pm$ 1.93
Category 67.97 $\pm$ 1.82 67.04 $\pm$ 1.89 65.30 $\pm$ 1.76 63.28 $\pm$ 1.92
Table 3: Classification results with binary and scaled labels on simple and category styled prompts, in which the scales were true-only generations taught by GPT-4o-mini.
@@
Original Augmented (Extended)
@@
[…]
Table 4: Classification results with and without newly augmented non-hateful meme data. Note: Until specified, we have used the base variant of the models mentioned, and the size is written as an approximate number of parameters in the model.
@@
We present the results and implications of our data augmentation strategies, evaluated via prompting and fine-tuning. While we report both accuracy and weighted F1-score, we focus on F1-score due to its greater sensitivity in class-imbalanced classification tasks.
@@
The impact of prompt styling varies significantly depending on the task complexity (binary vs. scaled). Under the binary label setting, the simple prompt consistently outperforms the category prompt across unimodal models. For instance, RoBERTa achieves an F1-score of 59.13 with the simple prompt compared to 52.87 with the category prompt. This suggests that for the simpler binary classification task, adding categorical definitions may introduce unnecessary noise to the inputs of smaller models.
However, for the augmented (scaled) dataset, this trend reverses. Both BERT and RoBERTa achieve higher F1-scores using the category prompt (e.g., RoBERTa improves from 57.42 to 58.24). Similarly, for Multimodal prompting with InternVL2, the category prompt provides a substantial boost on scaled labels, raising the F1-score from 60.00 to 64.74. This indicates that as the label space becomes more nuanced (scaled), providing categorical context via prompts becomes beneficial, helping models discern subtle degrees of hatefulness.
@@
Multimodal fine-tuning with InternVL2 achieves the highest performance on the binary dataset, with a peak F1-score of 67.28 (simple), significantly outperforming the prompting-only setup (F1: 54.51). This confirms the advantage of full model adaptation for the standard binary classification task.
Conversely, on the scaled dataset, multimodal prompting demonstrates surprising efficacy. The InternVL2 model using the category prompt achieves an F1-score of 64.74, which is competitive with and notably slightly higher than the best fine-tuning result on scaled labels (F1: 63.28). This suggests that for complex, scalar outputs, a well-structured prompt may leverage the inherent knowledge of large multimodal models as effectively as fine-tuning on a smaller augmented dataset.
@@
  • Task complexity dictates prompt utility: While simple prompts suffice for binary tasks, structured (category) prompts are crucial for optimizing performance on nuanced, scaled annotations across both small and large models.
  • Prompting competes with Fine-tuning on Scaled Data: On the augmented dataset, prompting InternVL2 with category definitions yields results comparable to fine-tuning (F1 64.74 vs 63.28), highlighting the potential of zero-shot prompting for granular hatefulness detection without expensive training.
  • Unimodal Robustness: Unlike previous iterations, standard unimodal models (BERT, RoBERTa) show they can effectively utilize categorical context to improve performance when the target labels are scaled, rather than binary.
@@
To assess the value of augmenting our training set with counterfactually neutral memes, we fine-tune a diverse range of popular unimodal and multimodal models, including BERT (Devlin et al. 2019), RoBERTa (Liu et al. 2019), ALBERT (Lan et al. 2020), ViT (Wu et al. 2020), BEiT (Bao et al. 2022), DeiT (Touvron et al. 2021), and CLIP (Radford et al. 2021), using identical hyperparameters across both the original and extended datasets. We also run zero-shot and fine-tuning experiments on SmolVLM (Marafioti et al. 2025), and Qwen-VL (Bai et al. 2025) series. We aim to determine whether exposure to visually and lexically similar non-hateful memes improves model generalization and robustness (Table 4).
@@
Misclassification has asymmetric costs: false negatives risk under-moderation of harmful content, while false positives risk over-moderation. Our evaluation therefore, prioritizes weighted F1 in addition to accuracy. Across most model types, the extended dataset yields modest gains in accuracy (typically <<1%), but more notable improvements in F1-score, indicating better handling of class imbalance and edge cases. For example:
  • Text Unimodal: BERT improves from 55.50 to 57.95 in F1 (+2.45), with a negligible drop in accuracy (59.08 to 58.20). RoBERTa shows a larger relative F1 gain, from 47.18 to 52.54 (+5.36), despite a slight accuracy decrease. ALBERT also benefits significantly, jumping from 54.96 to 58.54 F1.
  • Vision Unimodal: Vision-only models generally lag behind text models. BEiT shows the best improvement (F1: 53.38 to 55.43), whereas DeiT actually suffers a slight performance drop (F1: 52.58 to 51.99) on the augmented data, suggesting it may struggle to resolve the visual ambiguity introduced by counterfactual examples.
  • Multimodal Fine-tuning: Augmentation proves highly effective for VLM fine-tuning. SmolVLM improves from 51.12 to 54.53 F1. The state-of-the-art Qwen2-VL achieves the highest overall performance, rising from 69.59 to 71.31 F1, demonstrating that larger model capacities can better leverage the nuanced distinctions in the augmented data.
Zero-shot vs. Fine-tuning Capabilities
We further evaluate the zero-shot capabilities of recent Vision-Language Models (VLMs) to understand their inherent understanding of hateful memes without task-specific training (Table 5). We observe that model scale plays a critical role in zero-shot performance.
  • Small Model Gap: The smaller SmolVLM (2B) struggles in the zero-shot setting (F1: 43.59), significantly underperforming its fine-tuned counterpart (F1: 54.53). This highlights that smaller VLMs rely heavily on fine-tuning to adapt to the specific distribution of hateful memes.
  • Large Model Generalization: In contrast, Qwen2.5-VL (3B) and Qwen2-VL (7B) demonstrate remarkable zero-shot robustness. The 7B model achieves 69.13 F1 zero-shot, which is comparable to the 71.31 F1 achieved after fine-tuning on the augmented dataset.
  • Surprising Efficiency: Notably, the zero-shot Qwen2.5-VL (3B) achieves an F1 of 63.79, which actually outperforms the fine-tuned Qwen2.5-VL on the augmented dataset (62.01). This suggests that for certain architectures, the pre-trained knowledge base may be more robust than a model fine-tuned on a relatively small, domain-specific dataset.
Category
Model Size Accuracy F1-score
Multimodal (zeroshot ICL) SmolVLM 2B 58.68 $\pm$ 1.84 43.59 $\pm$ 2.23
Qwen2.5-VL 3B 63.52 $\pm$ 1.77 63.79 $\pm$ 1.75
Qwen2-VL 7B 69.82 $\pm$ 1.77 69.13 $\pm$ 1.85
Table 5: Classification results with different VLMs in zero-shot setting.
Implications
These findings corroborate that counterfactual data augmentation enhances model robustness, especially in F1, by reducing overfitting to superficial features and encouraging deeper semantic reasoning. Such augmentation strategies are particularly effective when models have sufficient capacity to leverage subtle distinctions between hateful and non-hateful content.
@@
Example 1 Example 2 Example 3 Example 4
Examples
HS*
0 5 3 4
Example 1 Example 2 Example 3 Example 4
Hateful Meme
Non-hateful Augmentation
@@
To evaluate the reliability of the label augmentation strategy, we perform a human evaluation on a random sample (N=200). Evaluators were asked to judge whether the generated scales align with their own assessments. The average agreement rate is 92.5%, indicating strong reliability of the machine annotations. As shown in Table 6, most disagreements arise from indirect or context-dependent hate speech, which can be subtle and challenging even for advanced models such as GPT-4o-mini. In Example 1, the phrase “boneless jungle ham”, paired with watermelon imagery, constitutes a specific derogatory slur targeting African Americans. Example 2 is better characterized as edgy or historical humor rather than explicit hate speech: although it references the violent context of WWII atomic bombings, it alludes to a historical event rather than expressing hostility toward Japanese people. These cases underscore the difficulty of identifying implicit and contextual hate and point to the need for improved modeling of sociocultural nuance.